### v1

# arXiv:2603.00276v1[math.OA] 27 Feb 2026

Abstract
If $G$ is a compact group, continuous normalized positive definite functions are in one-to-one correspondence with unital quantum channels acting as Fourier multipliers on the group von Neumann algebra $\mathrm{VN}(G)$. We study the convex geometry of the convex set $\mathrm{P}_{1}(G)$ of normalized positive definite functions, equipped with the topology induced by the norm topology of the Fourier algebra $\mathrm{A}(G)$, and its relation with the structure of $\mathrm{VN}(G)$. We show that the von Neumann algebras of two compact groups $G$ and $H$ are *-isomorphic if and only if the convex sets $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$ are affinely homeomorphic. We also describe the group of affine homeomorphisms of $\mathrm{P}_{1}(G)$ in terms of Jordan *-automorphisms of $\mathrm{VN}(G)$.
00footnotetext: 2020 Mathematics subject classification: 94A40, 81P45.
Key words: group von Neumann algebra, positive definite function, convex set, affine homeomorphism, Fourier multiplier quantum channel.
Contents
1 Introduction
Quantum channels are the basic morphisms of quantum information theory. In the Heisenberg picture, a quantum channel is a weak* continuous unital completely positive map between von Neumann algebras. In this paper, we focus on a natural subclass of quantum channels arising from harmonic analysis, namely those channels which act as Fourier multipliers on group von Neumann algebras.
[…]
Given a continuous positive definite function $\varphi\colon G\to\mathbb{C}$ with $\varphi(e)=1$, one can define, according to [KaL18, Proposition 5.4.9, p. 184], a weak* continuous unital completely positive map $M_{\varphi}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$ defined by
$M_{\varphi}(\lambda_{s})=\varphi(s)\lambda_{s},\quad s\in G.$
Such maps are the Fourier multiplier quantum channels associated with $G$. They form the core of the noncommutative harmonic analysis approach to quantum information. In [Arh25], the first author investigated Fourier multiplier channels, with a particular emphasis on finite groups, and an exhaustive description is given for concrete examples, such as the quaternion group $\mathbb{Q}_{8}$ and the dihedral group $\mathbb{D}_{4}$. As a byproduct, it is shown, remarkably, that each quantum channel acting on the matrix algebra $\mathrm{M}_{2}$ is a part of a Fourier multiplier on one of these groups.
Guided by the motivation to understand the differences between the structure of the spaces of Fourier multipliers on groups such as $\mathbb{Q}_{8}$ and $\mathbb{D}_{4}$, the aim of the present paper is to study the convex geometry of the set of Fourier multiplier channels through the convex geometry of positive definite functions when $G$ is compact. Let $\mathrm{P}_{1}(G)$ be the set of continuous positive definite functions $\varphi$ on $G$ such that $\varphi(e)=1$. The bijective correspondence $\varphi\mapsto M_{\varphi}$ is affine, and it turns the study of convex decompositions of Fourier multiplier channels into the study of convex decompositions inside $\mathrm{P}_{1}(G)$.
Our first observation is that the relevant convex set is more naturally described in terms of normal states of $\mathrm{VN}(G)$. Using Eymard’s identification $\mathrm{A}(G)\cong\mathrm{VN}(G)_{*}$, we show that $\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$ can be identified affinely and homeomorphically with the normal state space of $\mathrm{VN}(G)$. This implies that a *-isomorphism $\mathrm{VN}(G)\cong\mathrm{VN}(H)$ induces an affine homeomorphism between $\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$ and $\mathrm{A}(H)\cap\mathrm{P}_{1}(H)$.
We then concentrate on compact groups. In this case, we have $\mathrm{P}_{1}(G)=\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$. We equip $\mathrm{P}_{1}(G)$ with the induced topology coming from the norm topology of $\mathrm{A}(G)$, which identifies $\mathrm{P}_{1}(G)$ with the normal state space of $\mathrm{VN}(G)$. This allows us to use the general theory of facial structures of state spaces of operator algebras, in particular the correspondence between central projections and split faces developed in [AlS76], [AlS01] and [Ped18]. Our main result states that for compact groups $G$ and $H$, the von Neumann algebras $\mathrm{VN}(G)$ and $\mathrm{VN}(H)$ are *-isomorphic if and only if the convex sets $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$ are affinely homeomorphic.
The present approach is related in spirit to the work [LWW22], where norm one positive definite functions are used to extract structural information about finite groups. Our results provide at the same time a generalization and a characterization, completing the result [LWW22, Theorem 4.6 p. 99]. Finally, in the last section, we investigate the group of affine homeomorphisms of the set $\mathrm{P}_{1}(G)$ of normalized positive definite functions in the case where $G$ is compact. We give a very concrete description, via normal Jordan *-automorphisms of $\mathrm{VN}(G)$.
Structure of the paper.
In Section 2, we recall basic notions from convex geometry and the correspondence between projections in a von Neumann algebra and faces of its normal state space. In Section 3, we identify the set $\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$ with the normal state space of $\mathrm{VN}(G)$ and we explain how *-isomorphisms of group von Neumann algebras induce affine homeomorphisms at the level of positive definite functions. In Section 4, we specialize to compact groups and show how the convex structure of $\mathrm{P}_{1}(G)$ determines the group von Neumann algebra $\mathrm{VN}(G)$. Finally, in Section 5, we describe the group of affine homeomorphisms of the set $\mathrm{P}_{1}(G)$ of normalized positive definite functions in the case where $G$ is compact.
2 Preliminaries
Geometry
A point $x\in C$ is said to be extreme if it cannot be written in a nontrivial way as a convex combination of points of $C$, i.e., if for any $t\in(0,1)$ and $y,z\in C$ the equality
(2.1) $x=ty+(1-t)z\quad\text{implies}\quad x=y=z.$
Following [AlS01, p. 1], we say that a convex subset $F$ of a convex set $C$ is a face of $C$ if the following implication holds for any $x,y\in C$ and any $t\in]0,1[$
(2.2) $tx+(1-t)y\in F\quad\text{implies}\quad x,y\in F.$
Two convex subsets $F$ and $G$ of a real vector space $X$ are said to be affinely independent if every point $z$ in their convex hull $\mathrm{co}(F\cup G)$ can be uniquely expressed as a convex combination
z=tx+(1t)yz=tx+(1-t)y
where $0<t<1$, $x\in F$ and $y\in G$. Here ¡¡uniqueness¿¿ means uniqueness up to the obvious indeterminacy of $x$ when $t=0$ and of $y$ when $t=1$.
We say that a convex set $C$ is the free convex sum of two convex subsets $F$ and $G$ if $C=\mathrm{co}(F\cup G)$ and if $F$ and $G$ are affinely independent. In this case, we write
$C=F\oplus_{c}G.$
It is known that if $C=F\oplus_{c}G$ then the two sets $F$ and $G$ must be faces of $C$. We say that a face $F$ of $C$ is a split face if there exists another face $G$ such that $C=F\oplus_{c}G$. In this case, $G$ is unique. We call it the complementary split face of $F$.
States of operator algebras
We describe the correspondence between central projections and their split faces in the following result, which is based on [AlS01, Theorem 3.35 p. 143] and [AlS01, Proposition 3.40 p. 148]. See also [AlS03, Corollary 5.34 p. 162] and [AlS76, Theorem 11.5 p. 97].
Proposition 2.1
Let $\mathcal{M}$ be a von Neumann algebra with normal state space $\mathcal{K}$. We denote by $\mathcal{F}$ the set of all norm closed faces of $\mathcal{K}$ and by $\mathcal{P}$ the set of all orthogonal projections in $\mathcal{M}$, each equipped with its natural ordering. Then there exists a bijection $\Phi\colon\mathcal{P}\to\mathcal{F}$, $p\mapsto\mathrm{Face}(p)$, where
(2.3) $\mathrm{Face}(p)\overset{\mathrm{def}}{=}\{\omega\in\mathcal{K}:\langle p,\omega\rangle_{\mathcal{M},\mathcal{M}_{*}}=1\}.$
Moreover, this correspondence preserves the order, i.e.
(2.4) $p\leqslant q\Longleftrightarrow\mathrm{Face}(p)\subset\mathrm{Face}(q)$
and the disjointness, i.e.
(2.5) $pq=0\Longleftrightarrow\mathrm{Face}(p)\cap\mathrm{Face}(q)=\varnothing.$
Furthermore, $p$ is a central projection in $\mathcal{M}$ if and only if $\mathrm{Face}(p)$ is a split face of $\mathcal{K}$.
In this context, we say that $\mathrm{Face}(p)$ is the face supported by the projection $p$. Note that $\mathrm{Face}(0)=\emptyset$ and $\mathrm{Face}(1)=\mathcal{K}$.
Projections
First, we will use the following elementary result [Sun97, Proposition 2.5.4 p. 55].
Proposition 2.2
Suppose that $(H_{i})_{i\in I}$ is a family of closed subspaces of a Hilbert space $H$, which are pairwise orthogonal, with associated orthogonal projections $(p_{i})_{i\in I}$. Then the family $(p_{i})_{i\in I}$ is unconditionally summable, with respect to the strong operator topology. Moreover, the sum $\sum_{i\in I}p_{i}$ is the orthogonal projection on the closed subspace spanned by $\cup_{i\in I}H_{i}$.
Consider a (orthogonal) projection $p$ of a von Neumann algebra $\mathcal{M}$. Following [AlS03, Definition 3.28 p. 89], we say that $p$ is minimal if $p\not=0$ and if every projection $q$ of $\mathcal{M}$ such that $q\leqslant p$ is equal to 0 or to $p$. Equivalently, according to [KaR97b, Proposition 6.4.3 p. 420] $p$ is minimal if and only if we have the equality $p\mathcal{M}p=\mathbb{C}p$. By [KaR97b, p. 420], if $p$ belongs to the center $\mathrm{Z}(\mathcal{M})$ and is minimal in $\mathrm{Z}(\mathcal{M})$ then the reduced von Neumann algebra $\mathcal{M}p$ is a factor.
Example 2.3
Consider the von Neumann algebra $\mathrm{B}(H)$ of all bounded operators on a complex Hilbert space $H$. By [KaR97a, Example 5.1.7 p. 409], the minimal projections are the (orthogonal) projections on one-dimensional subspaces of $H$.
By [KaR97b, Corollary 6.5.3 p. 424], a factor is of type I if and only if it contains a minimal projection. Moreover, the same result states that a factor is of type $\mathrm{I}_{n}$ if and only if $1$ is the sum of $n$ minimal projections.
A von Neumann algebra $\mathcal{M}$ is said to be atomic if for every non-zero projection $p$ there exists a non-zero minimal projection $q$ such that $q\leqslant p$. The following characterization is [Hel09, Lemma 3.9 p. 93] (see also [Bla06, p. 354]).
Proposition 2.4
Let $\mathcal{M}$ be a von Neumann algebra with center $\mathrm{Z}(\mathcal{M})$. The following conditions are equivalent.
  1. 1.
    $\mathcal{M}$ is atomic.
  2. 2.
    There exists an orthogonal family $(p_{i})_{i\in I}$ of minimal projections of $\mathcal{M}$ such that $1=\sum_{i\in I}p_{i}$.
  3. 3.
    $\mathcal{M}$ is of type $\mathrm{I}$ and there exists an orthogonal family $(p_{i})_{i\in I}$ of minimal projections of $\mathrm{Z}(\mathcal{M})$ such that $1=\sum_{i\in I}p_{i}$.
  4. 4.
    $\mathcal{M}$ is a direct sum of type $\mathrm{I}$ factors.
Harmonic analysis
Let $G$ be a locally compact group. Recall [Dix77, p. 286] [KaL18, Definition 1.4.15 p. 22] that a function $f\colon G\to\mathbb{C}$ is said to be positive definite if
(2.6) $\sum_{j,k=1}^{n}\xi_{j}\overline{\xi_{k}}f(s_{k}^{-1}s_{j})\geqslant 0$
for any $\xi_{1},\ldots,\xi_{n}\in\mathbb{C}$ and any $s_{1},\ldots,s_{n}\in G$. Let $\mathrm{P}_{1}(G)$ be the set of continuous positive-definite functions $\varphi$ on $G$ such that $\varphi(e)=1$. The sum of two positive-definite continuous functions is positive-definite. If $\varphi$ is a positive definite function and if $\lambda\geqslant 0$, then $\lambda\varphi$ is also a positive definite function.
3 A description of the normal state space of a group von Neumann algebra
Recall that Eymard’s identification [Eym64] gives an isometric isomorphism $\Theta_{G}\colon\mathrm{A}(G)\to\mathrm{VN}(G)_{*}$ of Banach spaces, characterized by the property that, for every $\varphi\in\mathrm{A}(G)$, there exists a unique normal linear functional $\omega\in\mathrm{VN}(G)_{*}$ such that
(3.1) $\varphi(s)=\omega(\lambda_{s}^{*}),\qquad s\in G.$
This means that $\Theta_{G}(\varphi)=\omega$.
Proposition 3.1
Let $G$ be a locally compact group. The isomorphism $\Theta_{G}\colon\mathrm{A}(G)\to\mathrm{VN}(G)_{*}$ restricts to an affine homeomorphism
$\Theta_{G}\colon\mathrm{A}(G)\cap\mathrm{P}_{1}(G)\to\{\omega\in\mathrm{VN}(G)_{*}^{+}:\omega(1)=1\},$
which allows us to identify the subset $\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$ with the normal state space $\{\omega\in\mathrm{VN}(G)_{*}^{+}:\omega(1)=1\}$ of the von Neumann algebra $\mathrm{VN}(G)$.
Proof : Let $\omega\in\mathrm{VN}(G)_{*}^{+}$ with $\omega(1)=1$ and define the function $\varphi_{\omega}\colon G\to\mathbb{C}$ by $\varphi_{\omega}(s)\overset{\mathrm{def}}{=}\omega(\lambda_{s}^{*})$, where $s\in G$. For any finitely supported function $\xi\colon G\to\mathbb{C}$, we have
$\sum_{s,t\in G}\varphi_{\omega}(t^{-1}s)\overline{\xi_{t}}\xi_{s}=\sum_{s,t\in G}\omega(\lambda_{t^{-1}s}^{*})\overline{\xi_{t}}\xi_{s}=\sum_{s,t\in G}\omega(\lambda_{s}^{*}\lambda_{t})\overline{\xi_{t}}\xi_{s}$
$=\omega\Big[\Big(\sum_{s\in G}\xi_{s}\lambda_{s}^{*}\Big)\Big(\sum_{t\in G}\overline{\xi_{t}}\lambda_{t}\Big)\Big]=\omega\Big[\Big(\sum_{s\in G}\overline{\xi_{s}}\lambda_{s}\Big)^{*}\Big(\sum_{t\in G}\overline{\xi_{t}}\lambda_{t}\Big)\Big]\geqslant 0,$
since $\omega$ is positive. Hence $\varphi_{\omega}$ is positive definite. Moreover, we have $\varphi_{\omega}(e)=\omega(1)=1$. Since $\omega$ is normal, we have $\varphi_{\omega}\in\mathrm{A}(G)$ and $\Theta_{G}(\varphi_{\omega})=\omega$. Thus $\varphi_{\omega}\in\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$.
Conversely, let $\varphi\in\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$. Consider the linear form $\omega\overset{\mathrm{def}}{=}\Theta_{G}(\varphi)$ of $\mathrm{VN}(G)_{*}$. For any finite sum $a=\sum_{s}a_{s}\lambda_{s}$, we compute
$\omega(a^{*}a)=\omega\Big(\big(\sum_{s}a_{s}\lambda_{s}\big)^{*}\big(\sum_{t}a_{t}\lambda_{t}\big)\Big)=\sum_{s,t\in G}\overline{a_{s}}a_{t}\omega(\lambda_{s^{-1}t})$
[…]
Consequently, $\omega$ is positive on the *-algebra generated by $(\lambda_{s})_{s\in G}$, which is weak* dense in $\mathrm{VN}(G)$. We conclude that it is a positive normal functional on $\mathrm{VN}(G)$. Furthermore, observe that $\omega(1)=\varphi(e)=1$.
Since $\Theta_{G}$ is an isometry, the restriction of $\Theta_{G}\colon\mathrm{A}(G)\to\mathrm{VN}(G)_{*}$ is an affine homeomorphism for the norm topology of $\mathrm{A}(G)$ and the norm topology of $\mathrm{VN}(G)_{*}$.
Proposition 3.2
Let $G$ and $H$ be locally compact groups. Suppose that there exists a *-isomorphism $\mathrm{VN}(G)\cong\mathrm{VN}(H)$. Then the convex sets $\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$ and $\mathrm{A}(H)\cap\mathrm{P}_{1}(H)$ are affinely homeomorphic, where we use the induced topologies by the norm topologies of $\mathrm{A}(G)$ and $\mathrm{A}(H)$.
Proof : Let $\theta\colon\mathrm{VN}(G)\to\mathrm{VN}(H)$ be a *-isomorphism. Recall that by [Li92, Proposition 1.12.3 p. 61], the map $\theta$ is a weak* homeomorphism. Now, consider the preadjoint $\theta_{*}\colon\mathrm{VN}(H)_{*}\to\mathrm{VN}(G)_{*}$ defined by
(3.2) $\theta_{*}(\omega)=\omega\circ\theta,\quad\omega\in\mathrm{VN}(H)_{*}.$
It is linear, isometric, and maps $\mathrm{VN}(H)_{*}^{+}$ onto $\mathrm{VN}(G)_{*}^{+}$. Moreover, we have the equivalence
[…]
Hence $\theta_{*}$ maps the normal state space of $\mathrm{VN}(H)$ onto the normal state space of $\mathrm{VN}(G)$, affinely and homeomorphically. Therefore, the composition
$T\overset{\mathrm{def}}{=}\Theta_{G}^{-1}\circ\theta_{*}\circ\Theta_{H}\colon\mathrm{A}(H)\cap\mathrm{P}_{1}(H)\to\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$
is an affine homeomorphism.
Recall that by definition [KaL18, Definition 2.1.5 p. 40], the Fourier-Stieltjes algebra $\mathrm{B}(G)$ of a locally compact group $G$ consists of all finite linear combinations of continuous positive definite functions. It is known that $\mathrm{A}(G)\subset\mathrm{B}(G)$, see [Pis01, p. 54]. According to [Eym64, p. 209], if $G$ is compact then we have $\mathrm{B}(G)=\mathrm{A}(G)$. In particular case, we have
(3.3) $\mathrm{A}(G)\cap\mathrm{P}_{1}(G)=\mathrm{P}_{1}(G).$
From Proposition 3.2, we deduce the following result.
Proposition 3.3
Let $G$ and $H$ be compact groups. Suppose that there exists a *-isomorphism $\mathrm{VN}(G)\cong\mathrm{VN}(H)$. Then the convex sets $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$ are affinely homeomorphic, where we use the induced topologies by the norm topologies of $\mathrm{A}(G)$ and $\mathrm{A}(H)$.
4 Convex structure of $\mathrm{P}_{1}(G)$ and structure of the group von Neumann algebra
Topology on $\mathrm{P}_{1}(G)$.
Assume that $G$ is compact. Then every $\varphi\in\mathrm{P}_{1}(G)$ belongs to $\mathrm{A}(G)$ and satisfies $\left\|\varphi\right\|_{\mathrm{A}(G)}=\varphi(e)=1$. Hence $\mathrm{P}_{1}(G)=\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$ and, via Eymard’s identification $\mathrm{A}(G)\cong\mathrm{VN}(G)_{*}$, the set $\mathrm{P}_{1}(G)$ identifies affinely with the normal state space of $\mathrm{VN}(G)$. We equip $\mathrm{P}_{1}(G)$ with the induced topology coming from the norm topology of $\mathrm{A}(G)$ (equivalently, the norm topology of $\mathrm{VN}(G)_{*}$). When we speak of closed faces of $\mathrm{P}_{1}(G)$ in the sequel, we mean closed for this topology.
Facial structure of $\mathrm{P}_{1}(G)$ and von Neumann algebra $\mathrm{VN}(G)$
Recall that if $G$ is compact we have a *-isomorphism
(4.1) $\mathrm{VN}(G)=\prod_{\pi\in\operatorname{Irr}(G)}\mathrm{M}_{d_{\pi}},$
where $\operatorname{Irr}(G)$ is the set of all equivalence classes of irreducible unitary representations. Here $d_{\pi}$ is the degree of $\pi\in\operatorname{Irr}(G)$. We start with the following result.
Lemma 4.1
Let $G$ be a compact group. There exists a maximal family $(F_{i})_{i\in I}$ of mutually disjoint minimal split non-empty closed faces of the convex set $\mathrm{P}_{1}(G)$.
Proof : We justify the existence of the family by Zorn’s lemma, which states that if every totally ordered subset in a non-empty partially ordered set $P$ has an upper bound in $P$, then $P$ admits a maximal element. Let $\mathcal{S}$ be the set of all families $\mathcal{E}$ of mutually disjoint minimal split non-empty closed faces of $\mathrm{P}_{1}(G)$, partially ordered by inclusion.
Observe that the center of the matrix algebra $\mathrm{M}_{d_{\pi}}$ is $\mathbb{C}1$ by [KaR97a, Example 5.1.7 p. 309]. Note that the center
[…]
contains (non-zero) minimal projections, where minimality is understood relative to $\mathrm{Z}(\mathrm{VN}(G))$ rather than to $\mathrm{VN}(G)$.
Fix a non-zero minimal projection $p\in\mathrm{Z}(\mathrm{VN}(G))$. By Proposition 2.1, the set $\mathrm{Face}(p)$ is a split non-empty closed face of the convex set $\mathrm{P}_{1}(G)$, and it is minimal among split non-empty closed faces because $p$ is minimal among non-zero central projections, using the order preserving property (2.4) in Proposition 2.1. Indeed, if $F$ is a split non-empty closed face with $F\subset\mathrm{Face}(p)$, then $F=\mathrm{Face}(q)$ for some central projection $q$ and $\mathrm{Face}(q)\subset\mathrm{Face}(p)$ implies $q\leqslant p$ by (2.4). Since $p$ is minimal in $\mathrm{Z}(\mathrm{VN}(G))$, we have $q=0$ or $q=p$. As $F$ is non-empty, $q\neq 0$. Hence $q=p$ and $F=\mathrm{Face}(p)$. Therefore $\{\mathrm{Face}(p)\}\in\mathcal{S}$. So the set $\mathcal{S}$ is non-empty.
Now, let $(\mathcal{E}_{j})_{j\in J}$ be a totally ordered subset of $\mathcal{S}$. Set
$\mathcal{E}\overset{\mathrm{def}}{=}\bigcup_{j\in J}\mathcal{E}_{j}.$
Then $\mathcal{E}$ is again a family of mutually disjoint minimal split non-empty closed faces (disjointness and minimality are preserved because the chain is totally ordered by inclusion). Hence $\mathcal{E}$ is an upper bound of $(\mathcal{E}_{j})_{j\in J}$ in $\mathcal{S}$. By Zorn’s lemma, $\mathcal{S}$ admits a maximal element, that is, a maximal family $(F_{i})_{i\in I}$ of mutually disjoint minimal split non-empty closed faces of the set $\mathrm{P}_{1}(G)$.
Proposition 4.2
Let $G$ be a compact group. Given the convex set $\mathrm{P}_{1}(G)$ up to affine homeomorphism, we can recover the von Neumann algebra $\mathrm{VN}(G)$ up to *-isomorphism.
Proof : By Lemma 4.1, there exists a maximal family $(F_{i})_{i\in I}$ of mutually disjoint minimal split non-empty closed faces of the convex set $\mathrm{P}_{1}(G)$. By Proposition 2.1, for each $i\in I$, there exists a unique non-zero central projection $p_{i}\in\mathrm{VN}(G)$, which is minimal in $\mathrm{Z}(\mathrm{VN}(G))$, such that
[…]
Moreover, if $i\neq j$, the faces $F_{i}$ and $F_{j}$ are disjoint, hence $p_{i}p_{j}=0$ by the disjointness statement (2.5) in Proposition 2.1. Thus $(p_{i})_{i\in I}$ is a family of pairwise orthogonal minimal central projections.
Let $p\overset{\mathrm{def}}{=}\sum_{i\in I}p_{i}$, where the sum is taken in the strong operator topology. This sum exists because the projections $(p_{i})_{i\in I}$ are pairwise orthogonal, see Proposition 2.2.
Lemma 4.3
We have $p=1$.
Proof : Assume by contradiction that $p\neq 1$ and set $r\overset{\mathrm{def}}{=}1-p$. Then $r$ is a non-zero central projection. Since the von Neumann algebra $\mathrm{VN}(G)$ is atomic, its center $\mathrm{Z}(\mathrm{VN}(G))$ is atomic by Proposition 2.4, hence $r$ dominates a non-zero minimal central projection $q\leqslant r$. By Proposition 2.1, the face $\mathrm{Face}(q)$ is a non-empty split face, and it is minimal among non-empty split faces because $q$ is minimal and central. Furthermore, $q\leqslant r=1-p$ implies $qp_{i}=0$ for all $i\in I$. Hence $\mathrm{Face}(q)$ is disjoint from each face $F_{i}$ by Proposition 2.1. Therefore, the family $(F_{i})_{i\in I}\cup\{\mathrm{Face}(q)\}$ is a family of mutually disjoint minimal split non-empty faces, contradicting the maximality of $(F_{i})_{i\in I}$. Thus $r=0$, that is, $p=1$. In particular, we have
$\sum_{i\in I}p_{i}=1,$
with strong convergence.
Since it is a sum of matrix algebras, the von Neumann algebra $\mathrm{VN}(G)$ is of type $\mathrm{I}$ by [Li92, Proposition 6.7.2 p. 303] or Proposition 2.4. We deduce by [KaR97b, p. 422] that each reduced von Neumann algebra $\mathrm{VN}(G)p_{i}$ is also of type $\mathrm{I}$. Furthermore, since $p_{i}$ is minimal in $\mathrm{Z}(\mathrm{VN}(G))$ the von Neumann algebra $\mathrm{VN}(G)p_{i}$ is a factor by [KaR97b, p. 420].
Recall that by [KaR97b, Theorem 6.6.1 p. 426] a type $\mathrm{I}$ factor is *-isomorphic to $\mathrm{B}(\mathcal{H})$ for some Hilbert space $\mathcal{H}$ with dimension $\dim\mathcal{H}$ belonging to $\{1,2,\ldots\}\cup\{\text{infinite cardinals}\}$. So we have a *-isomorphism $\mathrm{VN}(G)p_{i}\cong\mathrm{B}(\mathcal{H}_{i})$ with $\dim\mathcal{H}_{i}=n_{i}$. Since we have the (unique) decomposition (4.1) into factors, every factor summand of $\mathrm{VN}(G)$ is finite-dimensional. Therefore, for each $i\in I$ there exists a unique $\pi_{i}\in\operatorname{Irr}(G)$ such that
$\mathrm{VN}(G)p_{i}\cong\mathrm{M}_{d_{\pi_{i}}}.$
In particular, $n_{i}=d_{\pi_{i}}$. Fix with [KaR97b, Corollary 6.5.3 p. 424] a maximal family $(p_{i,1},\ldots,p_{i,n_{i}})$ of mutually orthogonal minimal projections in $\mathrm{VN}(G)p_{i}$ such that
$p_{i}=\sum_{k=1}^{n_{i}}p_{i,k},$
where the sum is taken in the strong operator topology. For any integer $k\in\{1,\ldots,n_{i}\}$, set
$q_{i,k}\overset{\mathrm{def}}{=}p_{i,1}+\cdots+p_{i,k}.$
Then $(q_{i,k})_{k=1}^{n_{i}}$ is a strictly increasing family of projections in $\mathrm{VN}(G)p_{i}$, and by Proposition 2.1 we obtain a strictly increasing family of faces
[…]
We claim that (4.3) is maximal, and that $n_{i}$ is the maximal possible length of a strictly increasing chain of non-empty norm closed faces contained in $F_{i}$. Indeed, under a *-isomorphism $\mathrm{VN}(G)p_{i}\cong\mathrm{M}_{n_{i}}$, projections correspond to orthogonal projections onto subspaces of $\mathbb{C}^{n_{i}}$, hence each projection has a rank in $\{0,1,\ldots,n_{i}\}$. If
$e_{1}<e_{2}<\cdots<e_{m}$
is a strictly increasing chain of projections in $\mathrm{M}_{n_{i}}$, then
$\mathrm{rank}(e_{1})<\mathrm{rank}(e_{2})<\cdots<\mathrm{rank}(e_{m}),$
so necessarily $m\leqslant n_{i}$. Since (4.3) has length $n_{i}$, it is maximal. Consequently, the integer $n_{i}$ can be recovered from the facial structure of $F_{i}$ and is invariant under affine homeomorphism.
Moreover, the projections $(p_{i})_{i\in I}$ are pairwise orthogonal and satisfy $\sum_{i\in I}p_{i}=1$ strongly, hence the map $i\mapsto\pi_{i}$ is a bijection between $I$ and $\operatorname{Irr}(G)$. Thus, knowing the facial structure of $\mathrm{P}_{1}(G)$ allows us to recover the decomposition (4.1) up to *-isomorphism.
Corollary 4.4
Let $G$ and $H$ be compact groups. Suppose that there exists an affine homeomorphism between the convex sets $\mathrm{P}_{1}(G)\cong\mathrm{P}_{1}(H)$, where we use the induced topologies by the norm topologies of $\mathrm{A}(G)$ and $\mathrm{A}(H)$. Then the von Neumann algebras $\mathrm{VN}(G)$ and $\mathrm{VN}(H)$ are *-isomorphic.
Proof : If there is an affine homeomorphism between $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$, then they have the same facial structure. Now, it suffices to use Proposition 4.2.
A combination of Corollary 4.4 and Proposition 3.3 yields the following characterization.
Theorem 4.5
Let $G$ and $H$ be compact groups. The von Neumann algebras $\mathrm{VN}(G)$ and $\mathrm{VN}(H)$ are *-isomorphic if and only if the convex sets $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$ are affinely homeomorphic, where we use the induced topologies by the norm topologies of $\mathrm{A}(G)$ and $\mathrm{A}(H)$.
Remark 4.6
Let $G$ be a locally compact group. Recall that by [BHV08, Theorem C.5.2 p. 358], a normalized definite function $\varphi$ is an extremal point of the convex set $\mathrm{P}_{1}(G)$ if and only if the associated unitary representation $\pi_{\varphi}$ is irreducible.
5 Affine homeomorphisms of the set $\mathrm{P}_{1}(G)$ of normalized positive definite functions
Here, we describe the group of affine homeomorphisms of the set $\mathrm{P}_{1}(G)$ of normalized positive definite functions if the group $G$ is compact.
Jordan *-homomorphisms
A Jordan *-homomorphism [Sto13, p. 8] between von Neumann algebras $\mathcal{M}$ and $\mathcal{N}$ is a linear map $J\colon\mathcal{M}\to\mathcal{N}$ such that
$J(x^{*})=J(x)^{*}\quad\text{and}\quad J(xy+yx)=J(x)J(y)+J(y)J(x),\quad x,y\in\mathcal{M}.$
This means in particular that $J$ preserves the Jordan product $x\circ y\overset{\mathrm{def}}{=}\frac{1}{2}(xy+yx)$. It is worth noting that *-homomorphisms and anti-*-homomorphisms are Jordan homomorphisms. Moreover, by [Sto13, Theorem 3.2.4 p. 32], a normal *-Jordan automorphism $J\colon\mathrm{B}(H)\to\mathrm{B}(H)$ of $\mathrm{B}(H)$, where $H$ is a complex Hilbert space, is an *-automorphism or an anti-*-automorphism of $\mathrm{B}(H)$. In particular, for any $x\in\mathrm{B}(H)$ we have $J(x)=uxu^{*}$ or $J(x)=ux^{\top}u^{*}$ for some unitary $u\in\mathrm{B}(H)$.
The group of affine homeomorphisms of $\mathrm{P}_{1}(G)$
We will use the following result.
Lemma 5.1
Let $G$ be a compact group and write $\mathrm{VN}(G)=\oplus_{\pi\in\operatorname{Irr}(G)}\mathrm{M}_{d_{\pi}}$, with minimal central projections $(p_{\pi})_{\pi\in\operatorname{Irr}(G)}$. Fix $\pi\in\operatorname{Irr}(G)$. If $\mathcal{K}(\mathrm{VN}(G)p_{\pi})$ is the normal state space of $\mathrm{VN}(G)p_{\pi}$ then the map $\mathcal{K}(\mathrm{VN}(G)p_{\pi})\to\mathrm{Face}(p_{\pi})$, $\eta\mapsto\omega_{\eta}$, where $\omega_{\eta}\in\mathrm{VN}(G)_{*}$ is defined by
(5.1) $\omega_{\eta}(x)\overset{\mathrm{def}}{=}\eta(p_{\pi}xp_{\pi}),\qquad x\in\mathrm{VN}(G),$
is an affine homeomorphism. In particular, $\mathrm{Face}(p_{\pi})$ is affinely homeomorphic to the normal state space of $\mathrm{VN}(G)p_{\pi}\cong\mathrm{M}_{d_{\pi}}$.
Proof : Let $\eta\in\mathcal{K}(\mathrm{VN}(G)p_{\pi})$, that is, $\eta$ is a normal state on the von Neumann algebra $\mathrm{VN}(G)p_{\pi}$ with unit $p_{\pi}$. For any $x\in\mathrm{VN}(G)$, since $p_{\pi}$ is central we have $p_{\pi}xp_{\pi}=p_{\pi}x$, hence
[…]
By (2.3), we deduce that $\omega_{\eta}\in\mathrm{Face}(p_{\pi})$. Conversely, let $\omega\in\mathrm{Face}(p_{\pi})$, i.e. $\omega\in\mathrm{VN}(G)_{*}^{+}$ and $\omega(p_{\pi})=1$. Define the form $\eta_{\omega}$ on $\mathrm{VN}(G)p_{\pi}$ by
(5.2) $\eta_{\omega}(y)\overset{\mathrm{def}}{=}\omega(y),\qquad y\in\mathrm{VN}(G)p_{\pi}.$
Then $\eta_{\omega}$ is a normal positive functional on $\mathrm{VN}(G)p_{\pi}$ and $\eta_{\omega}(p_{\pi})=\omega(p_{\pi})=1$. So $\eta_{\omega}\in\mathcal{K}(\mathrm{VN}(G)p_{\pi})$.
The two constructions are inverse to each other: if $\eta\in\mathcal{K}(\mathrm{VN}(G)p_{\pi})$ and $y\in\mathrm{VN}(G)p_{\pi}$, then $p_{\pi}yp_{\pi}=y$ and hence
[…]
If $\omega\in\mathrm{Face}(p_{\pi})$ and $x\in\mathrm{VN}(G)$, then
[…]
Writing $x=\sum_{\pi}x_{\pi}$ according to the direct sum decomposition, we have $p_{\pi}xp_{\pi}=x_{\pi}\in\mathrm{VN}(G)p_{\pi}$. Moreover, $\omega(p_{\rho})=0$ for $\rho\neq\pi$ since $p_{\rho}\leqslant 1-p_{\pi}$ and
$0\leqslant\omega(p_{\rho})\leqslant\omega(1-p_{\pi})=1-\omega(p_{\pi})=0.$
Hence $\omega(x)=\omega(x_{\pi})=\omega(p_{\pi}xp_{\pi})$. So $\omega_{\eta_{\omega}}=\omega$.
Finally, both maps are clearly norm continuous, hence they are affine homeomorphisms.
If $G$ is a locally compact group, recall that we have an isometric isomorphism $\Theta_{G}\colon\mathrm{A}(G)\to\mathrm{VN}(G)_{*}$ of Banach spaces.
Proposition 5.2
Let $G$ be a compact group. Equip $\mathrm{P}_{1}(G)$ with the topology induced by the norm topology of $\mathrm{A}(G)$. Let $T\colon\mathrm{P}_{1}(G)\to\mathrm{P}_{1}(G)$ be an affine homeomorphism. Then there exists a normal Jordan *-automorphism $\theta\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$ such that
(5.3) $(T\varphi)(s)=\Theta_{G}(\varphi)\big(\theta^{-1}(\lambda_{s}^{*})\big),\quad\varphi\in\mathrm{P}_{1}(G),s\in G.$
Moreover, if we write the canonical decomposition $\mathrm{VN}(G)=\oplus_{\pi\in\operatorname{Irr}(G)}\mathrm{M}_{d_{\pi}}$ and denote by $p_{\pi}$ the minimal central projection supporting the summand $\mathrm{M}_{d_{\pi}}$, then there exists a permutation $\sigma$ of $\operatorname{Irr}(G)$ such that $d_{\sigma(\pi)}=d_{\pi}$ for all $\pi$, and for each $\pi$ there exists a unitary $u_{\pi}\in\mathrm{M}_{d_{\pi}}$ such that the restriction of $\theta$ to $\mathrm{VN}(G)p_{\pi}\cong\mathrm{M}_{d_{\pi}}$ is either
$x\mapsto u_{\pi}xu_{\pi}^{*},\quad\text{or}\quad x\mapsto u_{\pi}x^{\top}u_{\pi}^{*},\qquad x\in\mathrm{M}_{d_{\pi}},$
after identifying $\mathrm{VN}(G)p_{\sigma(\pi)}\cong\mathrm{M}_{d_{\pi}}$. Conversely, the construction (5.3) gives rise to an affine homeomorphism.
Proof : From (3.3) and Proposition 3.1, the set $\mathrm{P}_{1}(G)$ identifies affinely with the normal state space
$\mathcal{K}\overset{\mathrm{def}}{=}\{\omega\in\mathrm{VN}(G)_{*}^{+}:\omega(1)=1\}$
via the map $\Theta_{G}$. In particular, the given topology on $\mathrm{P}_{1}(G)$ agrees with the norm topology on the subspace $\mathcal{K}$ of $\mathrm{VN}(G)_{*}$. Transporting $T$ through $\Theta_{G}$, we obtain an affine homeomorphism
$\widetilde{T}\overset{\mathrm{def}}{=}\Theta_{G}\circ T\circ\Theta_{G}^{-1}\colon\mathcal{K}\to\mathcal{K}.$
Let $p\in\mathrm{Z}(\mathrm{VN}(G))$ be a central projection. By Proposition 2.1, the set
$\mathrm{Face}(p)=\{\omega\in\mathcal{K}:\langle p,\omega\rangle=1\}$
is a split norm closed face of $\mathcal{K}$, and every split norm closed face arises in this way. Since $\widetilde{T}$ is affine and bijective, it sends norm closed faces onto norm closed faces, and it preserves the property of being a split face. Hence there exists a unique central projection $\alpha(p)\in\mathrm{Z}(\mathrm{VN}(G))$ such that
$\widetilde{T}(\mathrm{Face}(p))=\mathrm{Face}(\alpha(p)).$
Using again Proposition 2.1, we see that $\alpha$ preserves inclusion of faces, hence preserves the order of central projections:
$p\leqslant q\iff\mathrm{Face}(p)\subset\mathrm{Face}(q)\iff\mathrm{Face}(\alpha(p))\subset\mathrm{Face}(\alpha(q))\iff\alpha(p)\leqslant\alpha(q).$
Similarly, $\widetilde{T}$ preserves disjointness of faces, hence $\alpha$ preserves orthogonality of central projections. In particular, $\alpha$ sends minimal central projections to minimal central projections. Since
$\mathrm{Z}(\mathrm{VN}(G))=\oplus_{\pi\in\operatorname{Irr}(G)}\mathbb{C}\,p_{\pi},$
there exists a permutation $\sigma$ of $\operatorname{Irr}(G)$ such that $\alpha(p_{\pi})=p_{\sigma(\pi)}$.
By Lemma 5.1, $\mathrm{Face}(p_{\pi})$ is affinely homeomorphic to the normal state space of $\mathrm{VN}(G)p_{\pi}\cong\mathrm{M}_{d_{\pi}}$. The integer $d_{\pi}$ can be recovered as the maximal length of a strictly increasing chain of non-empty norm closed faces contained in $\mathrm{Face}(p_{\pi})$ (as in the proof of Proposition 4.2). Since $\widetilde{T}$ is a homeomorphism, it preserves chains of faces and their lengths, hence $d_{\sigma(\pi)}=d_{\pi}$. For each $\pi$, the restriction
$\widetilde{T}_{\pi}\overset{\mathrm{def}}{=}\widetilde{T}\big|_{\mathrm{Face}(p_{\pi})}\colon\mathrm{Face}(p_{\pi})\to\mathrm{Face}(p_{\sigma(\pi)})$
is an affine homeomorphism between the state spaces of two matrix algebras of the same size.
It is a standard fact [Emch09, Theorem 2 p. 155] that every affine automorphism of the state space of $\mathrm{M}_{n}$ is implemented by a Jordan *-automorphism of $\mathrm{M}_{n}$. By [Sto13, Theorem 3.2.4 p. 32], such a map is a *-automorphism or a *-antiautomorphism. This means there exists a unitary $u\in\mathrm{M}_{n}$ such that
$\rho\mapsto u\rho u^{*}\quad\text{or}\quad\rho\mapsto u\rho^{\top}u^{*}$
on density matrices, equivalently on states.
Applying this to $\widetilde{T}_{\pi}$, we obtain a unitary $u_{\pi}$ and one of the two previous possibilities, after identifying $\mathrm{VN}(G)p_{\pi}\cong\mathrm{M}_{d_{\pi}}$ and $\mathrm{VN}(G)p_{\sigma(\pi)}\cong\mathrm{M}_{d_{\pi}}$. Define $\theta\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$ by sending $\mathrm{VN}(G)p_{\pi}$ onto $\mathrm{VN}(G)p_{\sigma(\pi)}$ with the corresponding block map. Then $\theta$ is a Jordan *-automorphism of $\mathrm{VN}(G)$, and by construction we have
$\widetilde{T}(\omega)=\omega\circ\theta^{-1},\qquad\omega\in\mathcal{K}.$
Transporting this identity back through $\Theta_{G}$ gives (5.4). The converse part is obvious.
Remark 5.3
[…]
(5.4) $\Theta_{G}\circ T(\varphi)=\Theta_{G}(\varphi)\circ\theta^{-1},\qquad\varphi\in\mathrm{P}_{1}(G).$
Declaration of interest
None.
Competing interests
The authors declare that they have no competing interests.
Data availability
No data sets were generated during this study.
Acknowledgment
Part of this work was carried out during a visit of the first author to Nankai University (China) in 2025, supported by the National Natural Science Foundation of China (Grant No. 12571143). The first author gratefully acknowledges the hospitality of Nankai University.